THE CONVERGENCE RATE OF FIXED-WIDTH SEQUENTIAL CONFIDENCE INTERVALS FOR A PARAMETER OF

AN EXPONENTIAL DISTRIBUTION

KEIICHI HIROSE^{1}, EIICHI ISOGAI^{2} AND CHIKARA UNO^{1}

1Graduate School of Science and Technology, Niigata University, 8050 Ikarashi 2-cho, Niigata 950-21, Japan

2Department of Mathematics, Niigata University, 8050 Ikarashi 2-cho, Niigata 950-21, Japan

(Received February 26, 1996; revised June 10, 1996)

Abstract. In this paper we consider sequential fixed-width confidence interval estimation for a parameter θ = aμ + bσ with a and b being given constants when the location parameter μ and the scale parameter σ of the negative exponential distribution are unknown. We investigate the rate of convergence of the coverage probability for fixed-width sequential confidence intervals of θ.

Key words and phrases: Two-parameter exponential, sequential confidence intervals, coverage probability, convergence rate.